a boat can go 60 mph in still water . it takesas long to go 560 miles upstream as it does to go downstream 640 miles. how fast is the current?
Answer by stanbon(48568)

a boat can go 60 mph in still water . it takes as long to go 560 miles upstream as it does to go downstream 640 miles. how fast is the current?
-------------------
Let current speed be "c".
Upstream DATA:
Distance = 560 miles ; Rate = (60-c) mph ; time = 560/(60-c) hrs
---------------------------------
Downstream DATA:
Distance = 640 miles ; Rate = (60+c) mph ; time = 640/(60+c) hrs
-----------------------------------
EQUATION:
time = time
560/(60-c) = 640/(60+c)
7/(60-c) = 8/(60+c)
420+7c = 480-8c
15c = 60
c = 4 mph (current speed)
===============
Cheers,
Stan H.